Determine the order of the given differential equation; also state whether the equation is linear or nonlinear.

  1. t2(d2y/dt2)+t(dy/dt)+2y=sint
  2. (1+y2)(d2y/dt2)+t(dy/dt)+y=et
1
Expert's answer
2021-11-28T19:10:22-0500

Solution;

(1)

Given;

"t^2\frac{d^2y}{dt^2}+t\frac{dy}{dt}+2y=sint"

The highest derivative present in the give differential equation is "\frac{d^2y}{dt^2}"

Therefore the order is 2.

A differential equation is linear if the dependent variable y and its various derivatives have degree one or zero the equation.

Highest power of the highest order derivative is 1.

Therefore the degree is 1.

It's a linear equation.

(2)

Given;

"(1+y^2)\frac{d^2y}{t^2}+t(\frac{dy}{dt})+y=e^t"

Highest derivative is "\frac{d^2y}{dt^2}" .

Therefore order is 2

The equation has "y^2\frac{d^2y}{dt^2}" which is a non linear term.

Hence is a non linear differential equation.